Margin calculation (Revenue-Expense)/Revenue = Margin or (x-y)/x=Z. In my case, I know the desired Margin (z) and know the Expense (y). I would like to plug in various values for (z) and have it calculate the revenue required, solve for X.
Thanks.

Answer by stanbon(75887)

Margin calculation (Revenue-Expense)/Revenue = Margin or (x-y)/x=Z. In my case, I know the desired Margin (z) and know the Expense (y). I would like to plug in various values for (z) and have it calculate the revenue required, solve for X.
------------------
Multiply both sides by x to get:
x-y = zx
Get the x-terms on one side:
x-zx = y
Factor out the "x":
x(1-z) = y
Solve for "x":
x = y/(1-z)
==================
Cheers,
Stan H.